A block of mass m rests on a horizontal table, executing SHM along horizontal with amplitude A when block
just starts to slip then the period ( = coefficient of friction) is

Grade:12th pass

2 Answers

Arun
4 years ago
F restoring =  F friction
F friction  = Ц mg
F restoring =  mw²A
Restoring force on the block = mω2A = μmg.
∴ Acceleration in the block = μg
Frequency of oscillation,
v  =   1  / 2 √acceleration / Displacement
=  1/2 π √μg / A
Vikas TU
4 years ago
Dear student 
Maximum restoring force = Limiting frictin force 
mw^2A =  μ mg
w^A = μg
w = sqrt( μg/A)
w = 2pie/T 
T = 2piesqrt(A/ μg)
Hope this will help